Reports 1-1 of 1 Clear search Modify search
AdV-PSL (Laser Bench setup)
turconi, soulard - 21:05 Wednesday 14 October 2020 (49661) Print this report
Master alignment into the Slave

Today we realigned the ML into the SL with the closest mirrors to the slave. The alignment changed because the ML beam was realigned on the EOM last week (see log entry 49585).

We could improve the error signal by a factor 2, it is now 1.2 Vpp.

Comments to this report:
chiummo - 21:30 Wednesday 14 October 2020 (49663) Print this report

After the work on the PSL, the IMC has been relocked.

The length of the IMC was tuned in order to have an fmoderr signal < 5e-6 (dataDisplay units), then the resonance of the RFC was found by tuning the length of the IMC around this value and the RFC was locked. The stored value of the thermal correction to the master laser was checked to be compliant with the correction needed to keep the RFC on resonance while the IMC length was close to an optimal fmoderr.

INJ is left with BPC engaged, PSTAB engaged, IMC locked, RFC locked, AA in drift control for the night.


turconi - 12:00 Tuesday 20 October 2020 (49694) Print this report

During this activity on the slave laser, the master laser temperture was set to 24.50, that was a value in the middle of the temperature range where no mode hops occur.

Next day, after the IMC and the RFC were locked, the temperature displayed value was 24.00.


Search Help